OpenStudy (anonymous):

Let f(x) = 4 - x and g(x) = x + 4. (f*g)(x)=

OpenStudy (akshay_budhkar):

Same logic as the previous question. \[(f \times g)x=f(x) \times g(x)=(4-x)(x+4)\] btw in the last question we have to be careful that \[x \neq -4 \] as the function wont be defined :)

OpenStudy (anonymous):

Oh, I see. O.k. Thanks again =)

OpenStudy (akshay_budhkar):

:)